# Breaking Information Cocoons: A Hyperbolic Graph-LLM Framework for Exploration and Exploitation in Recommender Systems

### v2

@@ 4.2 Semantic Representations Generation
In this subsection, we introduce a paradigm for distilling useful information from raw text descriptions of items and users into dense semantic embeddings, which will be used in our alignment framework to derive the final representation. Inspired by prior research (Ma et al., 2024; Ren et al., 2024) that leverages the power of large language models (LLMs) to extract useful user/item descriptions from raw, noisy textual data, we adopt a similar framework to generate user/item profiles with rich textual descriptions: $\mathcal{P}_{i}=\textit{LLMs}(\mathcal{S}_{i},\mathcal{T}_{i})$, where $\mathcal{S}_{i}$ represents system prompts to the LLMs, $\mathcal{T}_{i}$ denotes raw textual information, and $\mathcal{P}_{i}$ is the resulting user/item profile, consisting of sentence-level descriptions.
While we’ve derived textual information for users/items, it cannot be directly fused with the embeddings generated from hyperbolic collaborative filtering due to discrepancies in embedding geometry. To integrate both types of knowledge, we encode the textual information in embeddings using pre-trained text encoder models: $e_{i}=\textit{Encoder}(\mathcal{P}_{i})$. While LLM embeddings could be an alternative, they often introduce higher computational overhead and may not be optimized for alignment with hyperbolic representations, making encoders a more practical choice.
@@ 4.3 Hyperbolic Alignment
We propose maximizing mutual information between hyperbolic embeddings (collaborative information) and semantic embeddings (textual information) to create unified embeddings that integrate both knowledge types. To achieve this, we introduce a hyperbolic alignment loss, which reduces noise in representations and improves tail item recommendations by aligning information in the same hyperbolic space. A theoretical advantage of hyperbolic alignment over Euclidean alignment is discussed in the next subsection.
Before calculating the alignment loss, two steps are necessary: 1) the semantic embeddings must be adjusted to match the dimensionality of the collaborative embeddings, and 2) they should be transformed into a common representation space. To accomplish this, we first apply a multi-layer perceptron (MLP) to the semantic embeddings to ensure dimensional alignment. We then project the semantic embeddings into hyperbolic space so they can align with the hyperbolic collaborative embeddings:
$s_{i}=\exp_{x}(e_{i}),\quad s^{\prime}_{i}=\textit{MLP}(s_{i})$ (13)
where $\exp_{x}$ is defined as in Equation 3, $e_{i}$ represents the semantic embeddings from the encoder, $s_{i}$ is the dimension-aligned semantic embedding, and $s^{\prime}_{i}$ denotes the final projected semantic embedding in hyperbolic space. The hyperbolic alignment loss is then calculated as:
$\ell_{align}(i)=d_{\mathcal{H}}^{2}(h_{i},s^{\prime}_{i})$ (14)
where $d_{\mathcal{H}}$ represents the hyperbolic distance, and $h_{i}$ is the hyperbolic embedding with collaborative information.
@@ 4.4 Gradient Analysis of Hyperbolic Alignment
To further explore the impact of aligning semantic embeddings within the hyperbolic model, we analyze the gradient behavior with respect to nodes of varying norms in Proposition 4.1, proof can be found in Appendix A. This provides insight into the hierarchical preservation properties of the hyperbolic space, which are absent in Euclidean space.
Proposition 4.1.
@@
This proposition highlights that hyperbolic space offers adaptive gradient magnitudes based on node norms. Nodes with large norms (fine-grained preferences, small positional changes cause significant distance changes) receive smaller gradient updates, ensuring precise distance adjustments and preserving local structures. Meanwhile, nodes with smaller norms (abstract preferences, positional changes have less impact on distances) adjust more, refining global relationships and effectively capturing the hierarchical structure.
@@ 4.5 Hierarchical Representation Structure
In this subsection, we design a mechanism that allows users to balance the exploration-exploitation trade-off by constructing a hierarchical representation structure based on the learned hyperbolic embeddings.
@@
To address this, we introduce a hyperbolic hierarchical clustering method based on hyperbolic embeddings to build the hierarchy tree. Specifically, we iteratively group embeddings into clusters and use the centroid of each cluster as the data point to be clustered in the upper layer, referred to as a pseudo-cluster node. However, the number of cluster nodes in each layer is not predefined and must be determined. Inspired by Dasgupta’s cost (Dasgupta, 2016), which encourages pushing edge cuts as far down the tree as possible, the optimal tree is required to be binary. Therefore, we fix the proportion of cluster nodes between layers as $k=2$. The details of the algorithm are presented in Algorithm 1.
Algorithm 1 Hyperbolic Hierarchical Clustering
  Input: Hyperbolic embeddings $X$, layer proportion $k$
  Set the maximum number of layers $L=\log_{k}(|X|)$
  Initialize the points of bottom layer $D_{L}=X$
  for $l=L$ to $1$ (bottom to top) do
     repeat
        for each point $d_{l,i}\in D_{l}$ do
           Assign $d_{i}$ to the cluster with the nearest centroid:
           $c_{l,i}=\arg\min_{\mu\in M_{l}}d_{\mathcal{H}}(d_{i},\mu)$
        end for
        for each cluster $c_{l,i}\in C_{l}$ do
           $\mu_{l,i}=m_{\mathcal{H}}(\{d\ |\ d\in C_{l,i}\})$
        end for
     until Centroids $M_{l}$ do not change (convergence)
     Initialize the points of the upper layer $D_{l-1}=M_{l}$
  end for
@@
Complexity Analysis. To demonstrate the scalability and efficiency of our approach, we analyze the complexity of constructing and applying the hierarchical representation structure. Let $N$ denote the total number of users and items, and let $k$ be the number of recommendations per user. The hierarchy tree is built using iterative $k$-means clustering in hyperbolic space, where cluster nodes are assigned only to real node embeddings for efficiency.
First, computing and storing pairwise distances requires $O(N^{2}/2)$ operations. Each $k$-means iteration has complexity $O(i)$, and with $O(\log N)$ hierarchical levels, the total complexity of constructing the hierarchy tree is:$O\left(\frac{N^{2}}{2}+i\log N\right)$. For recommendation retrieval, finding the top-$k$ recommendations involves nearest neighbor search in hyperbolic space, which runs in $O(k\log N)$. Adjusting recommendations by retrieving ancestor clusters at layer $l$ requires only tree traversal, which runs in $O(k)$.
Overall, our hierarchical framework ensures efficient scaling while maintaining flexibility in recommendation adjustments. This design enables HERec to dynamically balance exploration and exploitation while ensuring scalability for large-scale recommender systems.
@@ 5 EXPERIMENTS
Data Amazon-books Yelp Google-reviews
Table 3: Statistics of the experimental data.
Dataset #User #Item #Interactions Density
All H20(%) T80(%)
[…]
Table 4: Performance on H20 (head) and T80 (tail) items, with the top hyperbolic results highlighted in bold. $\Delta_{\mathcal{H}}(\%)$ represents the improvement over the second-best hyperbolic baseline, which is underlined.
@@
5.1 Experimental Settings
Datasets. To evaluate our proposed model, we utilize three widely-used public datasets: Amazon-books, which contains user purchase behaviors within the book category on Amazon; Yelp, which records customer ratings and reviews for restaurants on Yelp; and Google-reviews, which includes user reviews and business metadata from Google Maps. Detailed statistics, including the distribution of head and tail items, are provided in Table 3. The high proportion of tail users (T80) highlights the prevalence of long-tail distributions, supporting the assumption of a power-law distribution in real-world data.
Evaluation Metrics. We assess our recommended items using both utility and diversity metrics, demonstrating that our model achieves precise recommendations for users’ preference while enhancing diversity in the recommendation. For utility, we adopt two widely used ranking metrics: Recall and NDCG, which measure the effectiveness of the model. To evaluate diversity, we employ three metrics: Distance Diversity (Div) (Ziegler et al., 2005; Zhang & Hurley, 2008), Shannon Entropy (H) (Mehta et al., 2012), and Expected Popularity Complement (EPC) (Vargas & Castells, 2011).
$H=-\sum_{i\in\text{set}(\sum_{u}R_{u})}p(i)\log p(i)$ (17)
Baselines. We compare our model with state-of-the-art baselines in both Euclidean and Hyperbolic spaces:
  • LightGCN (He et al., 2020): A lightweight graph convolution neural network model for recommendation by removing redundant neural modules in graph convolution.
  • TAG-CF (Ju et al., 2024): A test-time augmentation framework that enhances recommendation by performing a single message-passing step only at inference time.
  • LightGCL (Cai et al., 2023): A contrastive learning model that uses singular value decomposition (SVD) for robust graph augmentation without data augmentation.
  • SimGCL (Yu et al., 2022): A contrastive learning model that replaces graph augmentations in contrastive learning by introducing random noise during initialization.
  • HCCF (Xia et al., 2022): A collaborative filtering framework that integrates hypergraph and contrastive learning, capturing both local and global user dependencies.
  • HGCF (Sun et al., 2021): A hyperbolic GNN model that combines hypergraph learning with collaborative filtering using hyperbolic margin ranking loss.
  • HICF (Yang et al., 2022a): A hyperbolic recommendation model that enhances head and tail item performance with hyperbolic geometry, improving utility for long-tail items.
Training Details. We set embedding dimensions to 50 and a training batch size of 1024. Embeddings in hyperbolic space are optimized using Riemannian SGD (RSGD) (Bécigneul & Ganea, 2018; Sun et al., 2021), while the MLP adapter is optimized with Adam (Kingma, 2014). Training incorporates early stopping technique based on Recall@10 with 10 patience steps. We generate user and item profiles using GPT-3.5-turbo and encode them with text-embedding-ada-002 (Neelakantan et al., 2022) for Amazon-books and Yelp, and BERT (Devlin, 2018) for Google-reviews. All experiments are conducted on NVIDIA A40/A100 GPUs with 40GB memory.
@@ 5.1 Experimental Settings
5.2 Performance Comparison
Overall Comparison. The performance comparison across all models is presented in Table 1. Our HERec consistently outperforms the baselines in both utility and diversity metrics, achieving up to $5.49\%$ improvement in utility metrics and $11.39\%$ increase in diversity metrics. Notably, most baselines struggle to achieve strong performance in both utility and diversity simultaneously. For instance, while HICF performs second-best in utility metrics on the Amazon-books dataset, it lacks the same advantage in diversity metrics, where several baselines achieve better results. This highlights the superiority of our model, which achieves state-of-the-art performance in both utility and diversity: a feat not previously attained by any baseline.
Baselines with Feature Enhancement. To ensure a fair comparison, we further enhance Euclidean baselines by integrating feature information. Specifically, we align semantic embeddings with the original Euclidean embeddings following the structure in Section 4.2, but in Euclidean space. The results are presented in Table 2, with enhanced baselines denoted by “+”. The data shows that our HERec still surpasses all baselines in both utility and diversity. Interestingly, while the feature-enhanced versions of the Euclidean baselines generally improve in utility metrics, they do not exhibit a stable improvement in diversity metrics. This aligns with our theoretical analysis in Section 4.4, which demonstrates the advantages of our alignment in the hyperbolic space.
Additionally, to validate the effectiveness of our alignment framework in capturing semantics, we designed a profile decoder module to reconstruct user and item profiles from the learned hyperbolic embeddings. The results demonstrate that profiles reconstructed from our model’s embeddings exhibit higher alignment with the originals, confirming the model’s ability to effectively capture underlying semantic meaning. Detailed methodology and experimental results are provided in Appendix B.
@@ 5.2 Performance Comparison
Figure 5: Layer-wise norms within the hierarchy structure.
@@ 5.3 Ablation Study
We perform an ablation study to examine the contributions of two key components in our model: hyperbolic margin ranking loss and semantic alignment loss. To maintain meaningful evaluation, we focus solely on utility metrics, as random recommendations may artificially inflate diversity without practical utility. The following notations are used for our ablation variants: “w/o M” denotes the model without hyperbolic margin ranking loss, “w/o S” represents the model without semantic alignment loss, and “w/o M & S” indicates that both margin ranking and semantic alignment are excluded, leaving only the BPR loss.
Figure 3 shows that our full model outperforms all ablation variants, highlighting its overall effectiveness. Models lacking both margin ranking and semantic alignment perform the worst, while removing either module results in decreased utility, emphasizing the effectiveness of integration.
@@ 5.4 Hierarchical Representation Structure Analysis
We evaluated the effectiveness of our hierarchical representation structure through experiments simulating varying user preferences. Using a temperature parameter $\tau=0.5$ to control recommendation randomness, we analyzed performance across various hierarchical layers. As shown in Figure 4, utility metrics increase while diversity metrics decrease with larger layers. This trade-off aligns with the model’s design: larger layers focus on utility and specific interests for exploitation, while smaller layers prioritize diversity for exploration. This structure provides users with a flexible mechanism to adjust recommendations from niche, relevant items to broader, exploratory sets based on their preferences. A case study is provided in Appendix C.
Additionally, to validate the hierarchical structure’s capability to capture varying degrees of group preferences, we analyzed the average norm of node embeddings across layers (Figure 5). The results reveal a consistent increase at larger layers (farther from the center in hyperbolic space). As previously mentioned, hyperbolic geometry ensures that popular items or exploratory users are positioned closer to the origin, while niche items or users are positioned toward the boundaries. This finding aligns seamlessly with the properties of hyperbolic space, validating the intended design of the hierarchical structure: higher-level (smaller-layer) nodes effectively encapsulate abstract and aggregated preferences, whereas lower-level (larger-layer) nodes focus on fine-grained, personalized preferences.
@@ 6 CONCLUSION
In this work, we present HERec, a hyperbolic graph-LLM framework that effectively balances exploration and exploitation in recommender systems. Our approach integrates semantic and collaborative information within hyperbolic space, leveraging the unique properties of hyperbolic geometry. Furthermore, we introduce a personalized mechanism for balancing the exploration-exploitation trade-off using a hierarchical representation structure, empowering users to adjust recommendation preferences flexibly. Extensive experiments demonstrate that our model achieves state-of-the-art performance, achieving up to a $5.49\%$ improvement in utility metrics and an $11.39\%$ increase in diversity metrics. Remarkably, HERec is the first framework to excel in both utility and diversity simultaneously, marking a significant advancement in the field.
ACKNOWLEGDEMENTS
This work was partially supported by the National Science Foundation (NSF) IIS Div Of Information & Intelligent Systems 2403317. We also gratefully acknowledge partial support from an Amazon Research Award, funding generously provided by Snap Research, and resources made available by the Yale Office of the Provost.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.

### v3

@@ 4.2 Semantic Representations Generation
We propose a paradigm to distill useful information from raw text descriptions of items and users into dense semantic embeddings for our alignment framework. Building on prior work [20, 23] that leverages large language models (LLMs) to extract meaningful user/item descriptions from noisy textual data, we generate rich user/item profiles as $\mathcal{P}_{i}=\textit{LLMs}(\mathcal{S}_{i},\mathcal{T}_{i})$, where $\mathcal{S}_{i}$ are system prompts, $\mathcal{T}_{i}$ is raw textual input, and $\mathcal{P}_{i}$ contains sentence-level descriptions.
To integrate these profiles with hyperbolic collaborative filtering embeddings, we encode the textual information into embeddings using pre-trained text encoders: $e_{i}=\textit{Encoder}(\mathcal{P}_{i})$. While LLM embeddings are an alternative, they often incur higher computational overhead and are less optimized for alignment with hyperbolic representations, making pre-trained encoders a more practical choice.
@@ 4.3 Hyperbolic Alignment
To maximize mutual information between hyperbolic embeddings (collaborative information) and semantic embeddings (textual information), we introduce a hyperbolic alignment framework to create unified representations that integrate both knowledge types. To achieve this, we develop a hyperbolic alignment loss, which reduces noise in representations and improves tail item recommendations by aligning information in the same hyperbolic space. A theoretical advantage of hyperbolic alignment over Euclidean alignment is discussed in the next subsection.
Before calculating the alignment loss, two steps are necessary: 1) the semantic embeddings must be adjusted to match the dimensionality of the collaborative embeddings, and 2) they should be transformed into a common representation space. To accomplish this, we first apply a multi-layer perceptron (MLP) to the semantic embeddings to ensure dimensional alignment. We then project the semantic embeddings into hyperbolic space so they can align with the hyperbolic collaborative embeddings. The hyperbolic alignment loss is calculated as:
$\ell_{align}(i)=d_{\mathcal{H}}^{2}(h_{i},s_{i})$ (13)
where $d_{\mathcal{H}}$ represents the hyperbolic distance, $h_{i}$ is the hyperbolic embedding with collaborative information and $s_{i}$ denotes the projected semantic embedding in hyperbolic space.
@@ 4.4 Gradient Analysis of Hyperbolic Alignment
To further explore the impact of aligning semantic embeddings within the hyperbolic model, we analyze the gradient behavior with respect to nodes of varying norms in Proposition 1, detailed proof with analysis can be found in Appendix A. This provides insight into the hierarchical preservation properties of the hyperbolic space, which are absent in Euclidean space.
Proposition 1.
@@
This proposition highlights that hyperbolic space offers adaptive gradient magnitudes based on node norms. Nodes with large norms (fine-grained preferences, small positional changes cause significant distance changes) receive smaller gradient updates, ensuring precise distance adjustments and preserving local structures. Meanwhile, nodes with smaller norms (abstract preferences, positional changes have less impact on distances) adjust more, refining global relationships and effectively capturing the hierarchical structure. This inherent ability of hyperbolic space to modulate learning intensity based on an embedding’s position within the hierarchy offers a distinct advantage over the fixed-update approach of Euclidean geometry, ultimately facilitating more faithful and structured representations.
@@ 4.5 Hierarchical Representation Structure
In this subsection, we design a mechanism that allows users to balance the exploration-exploitation trade-off by constructing a hierarchical representation structure based on the learned hyperbolic embeddings. A detailed complexity analysis is provided in Appendix C.2.
@@
To address this, we introduce a hyperbolic hierarchical clustering method based on hyperbolic embeddings to build the hierarchy tree. Specifically, we iteratively group embeddings into clusters and use the centroid of each cluster as the data point to be clustered in the upper layer, referred to as a pseudo-cluster node. However, the number of cluster nodes in each layer is not predefined and must be determined. Inspired by Dasgupta’s cost [7], which encourages pushing edge cuts as far down the tree as possible, the optimal tree is required to be binary. Therefore, we fix the proportion of cluster nodes between layers as $k=2$. Details of the algorithm are presented in Appendix C.1.
@@ 5 EXPERIMENTS
5.1 Experimental Settings
Datasets. To evaluate our proposed model, we utilize three widely-used public datasets: Amazon-books, which contains user purchase behaviors within the book category on Amazon; Yelp, which records customer ratings and reviews for restaurants on Yelp; and Google-reviews, which includes user reviews and business metadata from Google Maps. Detailed statistics, including the distribution of head and tail items, are provided in Appendix B. The high proportion of tail users (T80) highlights the prevalence of long-tail distributions, supporting the assumption of a power-law distribution in real-world data. More implementation details can be found in Appendix B.1.
Evaluation Metrics. We assess our recommended items using both utility and diversity metrics, demonstrating that our model achieves precise recommendations for users’ preference while enhancing diversity in the recommendation. For utility, we adopt two widely used ranking metrics: Recall and NDCG, which measure the effectiveness of the model. To evaluate diversity, we employ three metrics: Distance Diversity (Div) [49, 43], Shannon Entropy (H) [21], and Expected Popularity Complement (EPC) [28]. For more details, please refer to Appendix B.2.
Baselines. We compare our model with state-of-the-art Euclidean and Hyperbolic baselines:
  • LightGCN [11]: A lightweight graph convolution neural network model for recommendation by removing redundant neural modules in graph convolution.
  • TAG-CF [12]: A test-time augmentation framework that enhances recommendation by performing a single message-passing step only at inference time.
  • LightGCL [3]: A contrastive learning model that uses singular value decomposition (SVD) for robust graph augmentation without data augmentation.
  • SimGCL [42]: A contrastive learning model that replaces graph augmentations in contrastive learning by introducing random noise during initialization.
  • HCCF [35]: A collaborative filtering framework that integrates hypergraph and contrastive learning, capturing both local and global user dependencies.
  • HGCF [25]: A hyperbolic GNN model that combines hypergraph learning with collaborative filtering using hyperbolic margin ranking loss.
  • HICF [36]: A hyperbolic recommendation model that enhances head and tail item performance with hyperbolic geometry, improving utility for long-tail items.
5.2 Performance Comparison
Overall Comparison. The performance comparison across all models is presented in Table 1, with statistical significance testing in Appendix D.2. Our HERec consistently outperforms the baselines in both utility and diversity metrics, achieving up to $5.49\%$ improvement in utility metrics and $11.39\%$ increase in diversity metrics. Notably, most baselines struggle to achieve strong performance in both utility and diversity simultaneously. For instance, while HICF performs second-best in utility metrics on the Amazon-books dataset, it lacks the same advantage in diversity metrics, where several baselines achieve better results. This highlights the superiority of our model, which achieves state-of-the-art performance in both utility and diversity: a feat not previously attained by any baseline.
Baselines with Feature Enhancement. To ensure a fair comparison, we further enhance Euclidean baselines by integrating feature information. Specifically, we align semantic embeddings with the original Euclidean embeddings following the structure in Section 4.2, but in Euclidean space. The results are presented in Appendix D.1, with enhanced baselines denoted by “+”. The data shows that our HERec still surpasses all baselines in both utility and diversity. Interestingly, while the feature-enhanced versions of the Euclidean baselines generally improve in utility metrics, they do not exhibit a stable improvement in diversity metrics. This aligns with our theoretical analysis in Section 4.4, which demonstrates the advantages of our alignment in the hyperbolic space.
Profile Decoder. To validate the effectiveness of our alignment framework in capturing semantics, we designed a profile decoder module to reconstruct user and item profiles from the learned hyperbolic embeddings. The results demonstrate that profiles reconstructed from our model’s embeddings exhibit higher alignment with the originals, confirming the model’s ability to effectively capture underlying semantic meaning. Detailed methodology and experimental results are provided in Appendix E.
Table 2: Performance on H20 (head) and T80 (tail) items, with the top hyperbolic results highlighted in bold. $\Delta_{\mathcal{H}}(\%)$ represents the improvement over the second-best hyperbolic baseline (underlined).
@@ 5.2 Performance Comparison
Figure 3: Layer-wise norms.
@@ 5.3 Ablation Study
We perform an ablation study to examine the contributions of two key components in our model: hyperbolic margin ranking loss and semantic alignment loss. To maintain meaningful evaluation, we focus solely on utility metrics, as random recommendations may artificially inflate diversity without practical utility. The following notations are used for our ablation variants: “w/o M” denotes the model without hyperbolic margin ranking loss, “w/o S” represents the model without semantic alignment loss, and “w/o M & S” indicates that both margin ranking and semantic alignment are excluded, leaving only the BPR loss. Figure 3 shows that our full model outperforms all ablation variants, highlighting its overall effectiveness. Models lacking both margin ranking and semantic alignment perform the worst, while removing either module results in decreased utility, emphasizing the effectiveness of integration.
Recent work [40, 39] has proposed performing message passing directly in hyperbolic space as an alternative to using the tangent space. We implemented this approach in place of our original hyperbolic graph collaborative filtering method. As shown in Appendix D.3, this alternative achieves comparable performance, making it a viable alternative.
@@ 5.4 Hierarchical Representation Structure Analysis
We evaluated the effectiveness of our hierarchical representation structure through experiments simulating varying user preferences. Using a temperature parameter $\tau=0.5$ to control recommendation randomness, we analyzed performance across various hierarchical layers. As shown in Figure 4, utility metrics increase while diversity metrics decrease with larger layers. This trade-off aligns with the model’s design: larger layers focus on utility and specific interests for exploitation, while smaller layers prioritize diversity for exploration. This structure provides users with a flexible mechanism to adjust recommendations from niche, relevant items to broader, exploratory sets based on their preferences. A case study is provided in Appendix C.3.
To further validate the hierarchical structure’s capability to capture varying degrees of group preferences, we analyzed the average norm of node embeddings across layers (Figure 3). The results reveal a consistent increase at larger layers (farther from the center in hyperbolic space). As previously mentioned, hyperbolic geometry ensures that popular items or exploratory users are positioned closer to the origin, while niche items or users are positioned toward the boundaries. This observation aligns with the theoretical properties of hyperbolic space and supports our hierarchical design: higher-level (smaller-layer) nodes effectively encapsulate abstract and aggregated preferences, whereas lower-level (larger-layer) nodes focus on fine-grained, personalized preferences.
@@ 6 CONCLUSION
In this work, we present HERec, a hyperbolic graph-llm framework that effectively balances exploration and exploitation in recommender systems. Our approach integrates semantic and collaborative information within hyperbolic space, leveraging the unique properties of hyperbolic geometry, supported by rigorous mathematical proof. Furthermore, we introduce a personalized mechanism for balance the exploration-exploitation trade-off using a hierarchical representation structure, empowering users to adjust recommendation preferences flexibly. Extensive experiments demonstrate that HERec achieves state-of-the-art performance, delivering up to a $5.49\%$ improvement in utility metrics and an $11.39\%$ increase in diversity metrics. Remarkably, HERec is the first framework to excel in both utility and diversity simultaneously, marking a significant advancement in the field.